### v1

# Using Half-Precision for GNN Training

Abstract.
Recent trends in lower precision, e.g. half-precision floating point, training have shown improved system performance and reduced memory usage for Deep Learning while maintaining accuracy. However, current GNN systems cannot achieve such goals for GNN, as our analyses show that they massively underperform while showing abnormal accuracy when using half-precision. These systems suffer from value overflow issues due to lowered precision, under-utilization of hardware resources, and poor training performance. To mitigate this, we introduce HalfGNN, a half-precision based GNN system. HalfGNN proposes novel techniques: new vector operations for half-precision data types that improve data load and reduction performance, and discretized SpMM that overcomes the value overflow and natively provides workload balancing. Such techniques improve hardware utilization, reduce memory usage, and remove atomic writes. Evaluations show that HalfGNN achieves on average of 2.30×2.30$\times$2.30 × speedup in training time over DGL (float-based) for GAT, GCN, and GIN respectively while achieving similar accuracy, and saving 2.67$\times$ memory.
1. Introduction
Graph neural networks (GNN) have become ever more present in many deep learning applications  (xu2018powerful, ; courbariaux2015binaryconnect, ; han2020graph, ; krahmer2003graph, ; wang2021relational, ; wu2022graph, ; ying2018graph, ; zitnik2018modeling, ). This is because graphs can model different kinds of phenomena across various domains. It has been shown that GNNs offer better performances than other approaches (hamilton2017inductive, ; xu2018powerful, ) for graph datasets. GNN relies on two key sparse kernels: SpMM (Sparse matrix-matrix multiplication), and SDDMM (Sampled dense-dense matrix multiplication) and their many variants. These two kernels are very different from a GeMM (General Matrix Multiplication) kernel due to the unstructured sparsity pattern present in the graph dataset.
GPU memory and computation capability have increased in recent years. However, it is still limited compared to the requirements of training large models. So, recent DL systems have explored lower precision floating-point data types such as half precision (16-bit floating point) instead of float (32-bit floating point) to utilize GPU resources efficiently. The half data type can theoretically provide faster computation and lower memory consumption than float. PyTorch (pytorch2019, ), a popular DL framework, has introduced mixed-precision training to integrate half data types. DGL (dgl2019, ), a popular GNN library, does provide support for half, which internally uses the Cusparse library for SpMM kernel, and in-house SDDMM. However, as we explain next, GNN remains to be explored using half precision.
Training Runtime Efficiency. We plotted the SpMM runtime using Cusparse, which DGL uses internally, in Fig. 1(a) for different feature lengths on two widely used datasets in GNN literature. The dataset and machine configurations are discussed in  section 5. The results show that half-precision SpMM is significantly slower than single-precision float. Similarly, Fig. 1(b) shows that DGL SDDMM does not improve runtime when the half-precision is used. These results indicate that the performance characteristics of sparse kernels for half-precision are not yet fully explored.
The performance characteristic is further complicated due to the availability of more hardware choices, such as tensor core, a hardware unit within GPU that performs faster computation on half-precision. To apply tensor cores to sparse kernels, we need to convert the sparse matrix to dense tiles; however, a GNN dataset may not be made up of dense tiles due to its unstructured sparsity of graph datasets. Unfortunately, TC-GNN (wang2023tc, ), which postulates tensor cores can provide better SpMM performance, actually causes huge performance degradation as shown by numerous prior works (Graphpy, ; GNNBench, ).
Training Accuracy. Most importantly, Fig. 1(c) shows that DGL achieves significantly lower training accuracy when using half for Ogb-Product and Reddit datasets. It would be acceptable if there were small losses in training accuracy due to loss in precision. Unfortunately, the training loss becomes NaN(not a number) after a couple of iterations. The observation begs the question: whether the model or the underlying system design and integration is responsible for significantly lower accuracy?
(a) SpMM Runtime Comparison
(b) SDDMM Runtime Comparison
(c) Training Accuracy Comparison
Figure 1. DGL half-precision Analysis over Float-based DGL: (a) Half-precision SpMM is very slow, (b) Half-precision SDDMM is similar in runtime than float-based SDDMM, (c) Low accuracy for GCN and GIN in half-precision
Contribution. These observations motivate us to explore and understand the GNN accuracy and performance characteristics when using half-precision, resulting in the proposed HalfGNN, the first ever half-precision based GNN System that achieves accuracy equivalent to float-based GNN and achieve significant runtime performance. We emphasize that the CUDA core is still the best for the computation as the tensor core is not suited due to the unstructured sparsity present in the graph datasets. Hence borrowing the existing idea from float-based sparse kernels becomes relevant. Unfortunately, adopting prior techniques fails to bring better performance because of several issues: (a) data-load does not achieve full memory coalescing in GPU due to the usage of half-precision as each GPU warp can only issue 64 bytes of data-load and not 128 bytes; (b) half-based arithmetic can only achieve sub-optimal throughput thereby lowering the performance; (c) current float-based workload balancing techniques use atomic instruction for conflicting writes, which is very costly for half-precision based operation.
To this end, we propose the following contribution:
$\bullet$ Adopting Half2 Data-Type for a Strong Baseline for Half-Precision Sparse Kernels. We adopt half2 data-type, a native feature of GPUs that treats two half-precision numbers as a single entity for memory operation (data-load and data-store instructions) and computations, to achieve full coalesced memory access and better arithmetic throughput. Adopting is a challenge that we overcome using Edge Feature Mirroring, Feature Padding for Odd Vertex Features, and introducing sub-warps to impart better thread utilization. This establishes an ideal baseline for half-precision kernels which we analyze and propose following techniques as they still fail to provide the envisioned performance and accuracy.
$\bullet$ More Vector Types for Improving SDDMM Performance. We introduce grouping of $4$ and $8$ half values, called (half4 and half8) respectively to exploit the native float2 and float4 support of GPUs to increase data-load performance by being able to issue more feature load instructions before the implicit memory barrier is encountered in the inter-thread communication for SDDMM reduction.
$\bullet$ Overflow and Discretized Reduction For SpMM. We analyze the root cause of NaN training error in GNNs. A vertex with a large neighborhood size, common in graphs, can produce INF (infinity) due to value overflow in SpMM reduction as half-precision has a limited numerical range. This INF then leads to NaN generation due to follow-up operations such as softmax involving arithmetic on two INF. However, GNN models indeed offer built-in mechanisms to protect against overflow such as degree-norm scaling in GCN which scales the output of SpMM back to half-precision range. But, merely the presence of such a mechanism is insufficient unless explored systematically. This work studies many such mechanisms across different GNN models and embeds them into system design through discretized reduction where only a finite batch of dot products are reduced and scaled (normalized) before handling the next batch of dot products, thus protecting the output against overflow.
The discretization is naturally suited to provide workload balancing in SpMM as one is free to choose the batch size, thereby allocating an equal amount of computation cores to each discretization unit leading to workload balancing. We demonstrate an edge-parallel and a vertex-parallel SpMM to show the generality of the proposed discretization for workload balancing and propose non-atomic write through usage of staging buffer and a follow-up kernel.
$\bullet$ Fear of Overflow and Frequent Data-Conversion. Further study shows that a naive mix-precision system performs excessive data-type conversion of tensors (float to half and vice-versa) or invokes many float-based kernels thereby not providing much benefits in moving to mix-precision training. E.g., the PyTorch exponential operation works in float due to fear of overflow. This operation is used in the GAT attention mechanism, introducing half to float on its input tensor and float to half on the output tensor. However, these conversions are not needed due to the built-in properties of the GAT attention mechanism that guarantee no overflow. This paper studies many such operations to truly incorporate half-precision.
Evaluations show that HalfGNN achieves state-of-the-art accuracy performance, along with massive improvement in runtime performance for half-precision. Briefly, we reduce the runtime on average of 2.30×2.30$\times$2.30 × and required memory by 2.67×2.67$\times$2.67 × on average over DGL for GAT, GCN, and GIN respectively. Furthermore, HalfGNN achieved 7.12×7.12$\times$7.12 ×, and 22.89×22.89$\times$22.89 × speedup over Cusparse half SpMM, and DGL half SDDMM.
We applied some of the performance optimizations of HalfGNN to Huang et al.  (huang2021understanding, ) and achieved 1.79×1.79$\times$1.79 × speedup for SpMM, validating the generality of the performance optimizations.
The remainder of the paper is organized as follows. We present the necessary background in section 2 and analyze the current GNN systems for half-precision, discuss related works, and present an overview in section 3. Detailed design is presented in section 4, evaluations in section 5, and conclusion in section 6.
2. Background
GPU. A program written for GPU is called a kernel. Computation units in GPUs are known as Symmetric Multiprocessors(SM). SM provides a programmable hardware cache, called shared memory, which the kernel can explicitly use to write and read. Programmers create and configure Cooperative Thread Arrays (CTA) or thread-blocks that are mapped to SM by the GPU driver. In a CTA, threads are usually grouped into 32 threads, called a warp, that executes computation in SIMT (single instructions multiple threads). Kernels are generally designed to access memory in a contiguous manner for better data-load performance.
Figure 2. A sample graph, its storage formats, and illustration of SpMM and SDDMM kernels: the two colors represent different vertex features.
Graph Storage Format and Tensor Types Fig. 2 shows a sample graph G = {V, E} where V is the vertex set and E is the edge set. We use graph theory and linear algebra terminologies (e.g., row, columns, and non-zero elements (NNZ)) to describe the graph and features of vertices and edges. The two most commonly used graph formats are Compressed Sparse Row (CSR) and Coordinate Objects (COO). COO uses two arrays of rowID and ColID where each $(rowID,colID)$ pair denotes an edge or non-zero element (NZE). The CSR also uses two arrays for its format: offset array points to the start of the neighborhood of each vertex or row. The degree defines the neighbor size of each vertex.
We also classify tensors involved in kernels, as vertex-level or edge-level tensors. For a graph having $|V|$ vertices and $|E|$ edges, each vertex having $|F|$ features would have a feature matrix of shape $|V|\times|F|$ (vertex-level tensor). For edge features, we get $|E|\times 1$ tensor (edge-level tensor).
Half Precision and Vector Data Types. Half-precision floating point numbers imply 16-bit floating point numbers and hence support a lower range of absolute values: $2^{-14}$ to $(2-2^{-10}).2^{15}$ i.e. $6.1\times 10^{-5}$ to $6.5\times 10^{4}$, which is very small compared to the range of single-precision float data-type. Hence, any value greater than the upper bound overflows to INF, while smaller values underflow to zero. Modern GPUs have provided support for half-precision arithmetic.
Vector types allows grouping of various lengths of basic type for data-load and computation. For example, two half numbers could be packed in one half2 vector data type, whose size is 32 bits. GPUs support native arithmetic on the half2 data type. On the other hand, float2 and float4 pack two and four float data. However, these two vector types support native data-load, but not arithmetic.
2.1. Deep Learning Traning and Sparse Kernels
Deep learning training process has two parts: forward and backward computation. The forward calls many kernels from the input layer to the output layer deriving prediction values. They are compared with the ground truth to calculate the loss. Backward uses the loss and invokes kernels from the output layer to the input layer to derive gradients and update the model parameters. The backward computation accesses state-tensor– results that are produced in the forward computation.
For sparse matrix dense matrix multiplication (SpMM) ($Y\leftarrow A_{w}X$), its backward computation needs to include both the SpMM ($\delta X\leftarrow A_{w}^{T}\delta Y$) and sampled dense dense matrix multiplication (SDDMM) ($\delta W_{e}\leftarrow A\odot(\delta YX^{T})$), where $\delta$ represents gradients of the tensors. The paper deals with these 2 classes of sparse kernels.
SpMM is categorized into 2 types. SpMMve is the usual SpMM, shown in Fig. 2d, where the edges contains features, and hence $A_{w}$ implies a graph topology ($A$) plus an edge-level tensor ($W_{e}$). It is then multiplied to a dense tensor X of size $|V|\times|F|$. A special case is SpMMv where all edge features are treated as 1.0, hence we do not need to store an explicit $W_{e}$. In SDDMM, shown in Fig. 2e, X and Y (or $\delta Y$) are two vertex-level tensors of size $|V|\times|F|$ while $A$ is meant to produce $W_{e}$ only (features/gradients on edges).
SpMMv is the only sparse kernel that is used in GCN (kipf2016semi, ) and GIN (xu2018powerful, ). Attention-based GNNs, such as GAT (gat18iclr, ) would require both SpMMve and SDDMM.
Sparse kernels broadly have 3 key steps: a) Data Load brings graph and features from the GPU global memory to the GPU registers of the computation unit. b) Actual Computation performs parallel computations to compute dot product of features; c) Reduction and Data-Store first reduces the computed dot products followed by writing to the output (GPU) memory. SpMM and SDDMM vary in how they perform the reduction phase. SpMM reduction happens across neighborhood dimensions, needing inter-thread communication only when the neighborhood of a vertex is allocated more than one warp to process, as in workload-balanced SpMM. SDDMM performs reduction across feature dimensions, so reduction mostly needs inter-thread communication.
Vertex-parallel (bell2009implementing, ) allocates equal computation to each vertex to process its neighborhood. Works such as (gale2020sparse, ; featgraphsc20, ; gespmmsc2020, ; fu2022tlpgnn, ; wu2021seastar, ) uses vertex-parallel variants. Workload imbalance occurs when certain vertices have large neighborhoods compared to others. Edge-parallel systems (yang2018design, ; gnnone, ) divide edges equally among computation units leading to a fully workload-balanced design. They introduce the need for inter-thread communication in SpMM for reduction and atomic write for data-store. Both types of parallel systems dedicate feature-parallel methods (yang2018design, ) when loading the vertex-level features leading to coalesced memory access for the same.
3. Mixed-Precision GNN Training: Analysis, Related Works, and Overview
Micikevicius et al (micikevicius2017mixed, ) shows that half-precision is as good as a float for diverse kinds of models without any accuracy loss. For this to happen two things need to be ensured during training: 1) as many forward and backward kernel operations can be carried out in half-precision as possible and storing state tensors in half-precision, and 2) the model weight updates must be in float, which automatically happens through mixed-precision module of PyTorch. Currently, no GNN system has applied this successfully. DGL fails to train correctly even on mid-size datasets of Reddit and Ogb-product (Fig. 1(c)), while its half-precision SpMM and SDDMM have not shown any performance improvement. Next, we present root cause analysis for abnormal accuracy and other system design issues, while discussing related works.
3.1. Analysis
We analyze half-precision DGL in more detail and found that a) half-precision SpMM causes value overflow when invoked by mixed-precision training; b) fear of overflow in mixed-precision forces PyTorch kernels to be invoked using float data-type in some cases due to less understanding about system integration to mixed-precision training. A naively designed GNN system in such cases either would invoke float-based sparse kernels or would be performing additional tensor conversion (float to half-precision). DGL does the latter. c) Half-precision kernels do not follow the current state-of-the-practice well-established in dense kernel counterparts. However, we also analyze various float-based GNN systems to show they would face similar issues if extended to half-precision or even borrow the idea from their dense counterparts. We elaborate on these findings next.
3.1.1. Value Overflow
GCN and GIN use SpMMv whose reduction phase causes output to overflow if a vertex has more neighbors irrespective of the system design. Overflow condition implies that the value range is too small to hold the output in half-precision, resulting in INF (plus or minus infinity). Even modest-sized datasets such as Ogb-Product and Reddit, have many vertices with large enough neighbors. The INF then produces NaN in the follow-up operations, e.g., softmax layer, which involves operation on two INF.
We next show that both the GNN models that rely on SpMMv offer built-in mechanisms to guard against overflow but have not been considered.
GCN. Eq. 1 shows a simplified computation involved in GCN. where, ${\mathbf{D}}$ is the degree matrix of graph ${\mathbf{A}}$. $\mathbf{H}^{(l)}$ is the activation of the previous layer, and $\mathbf{W}^{(l)}$ is the weight matrix of linear layer. The linear layer transforms the activation of the previous layer, which then serves as a dense input matrix to the SpMMv. This also includes a degree-norm scaling. GCN supports 3 different kinds of degree-norm: 1) left: divide the input dense tensor of SpMMv with degree; 2) right: divide the output tensor of SpMMv with degree; and 3) both: divide the input and output dense tensors of SpMMv with the square root of degrees separately which is shown in Eq. 1.
(1) $\mathbf{H}^{(l+1)}=\sigma\left({\mathbf{D}}^{-\frac{1}{2}}{\mathbf{A}}{\mathbf{D}}^{-\frac{1}{2}}(\mathbf{H}^{(l)}\mathbf{W}^{(l)})\right)$
The degree-norm scaling adds stability to feature values. To illustrate, consider the most used degree-norm, right, which simply divides the output features of every vertex produced by SpMMv by its degree. Hence, it is the built-in mechanism that can bring the SpMMv output within the half-precision range. However, the current systems perform degree-norm division after the reduction is over, by then overflow has already happened thus it is unable to protect it.
Some prior float-based GNN works have fused the degree-norm with SpMMv. However, our analysis shows that such approaches are not helpful because they also do degree-norm scaling after the reduction, hence they cannot protect against overflow in half-precision computation. Kindly note that if one uses the left degree-norm, there will be no overflow. However, during backward computation, the degree-norm happens after SpMMv where it is likely to overflow.
GIN. GIN has many variations, we focus on two variations. a) In the most common version, there is no stabilization operation like the degree-norm discussed for GCN. Therefore, there is no method to protect against the value overflow if one uses half-precision. b) DGL offers a few additional reduction operations in addition to the ‘sum’ for the SpMM operation in GIN. One of those options is ‘mean’ which is the same as SpMM followed by degree-norm as discussed in GCN. However, this option exists only in DGL and is implemented the same way as GCN, i.e., the degree-norm is called after SpMM for forward computation. Consequently, this version of GIN is susceptible to the same overflow issue as GCN, while GIN implemented by other float-based GNN systems would always overflow.
3.1.2. Fear of Overflow and Frequent Data-Conversion
This sub-section shows that integration of Custom Kernel System to Mixed-Precision Training has not been clearly understood, thereby half-precision sparse kernels are not invoked as we explain next with a specific example of attention-based GNN, such as GAT, followed by generalizing the observation.
Exponential Operation in GNNs. The final edge-level features in attention-based GNN, also called attention score ($\alpha_{ij}$) are derived using an edge-softmax operator as shown in Eq. 2. Here, the exponential of $e_{ij}$, called attention coefficients, is normalized with the cumulative sum of the exponential of all neighboring $e_{ij}$. $e_{ij}$ gives importance of neighbor $j$’s feature to vertex $i$. These are computed using a series of variants of SDDMM and SpMM kernels. However, PyTorch’s mixed-precision automatically invokes float-based exp operations instead of half-precision computation in Eq. 2.
(2) $\alpha_{ij}=\frac{\exp(e_{ij})}{\sum_{k\in\mathcal{N}_{i}}\exp(e_{ik})}.$
The exponential operation is immediately followed by an SDDMM variant to implement the division operation of Eq. 2 followed by SpMMve. All these follow-up sparse operations receive float type tensors due to this upgrade of just one operation (exponential) to float type. Hence, GNN systems would either a) perform data conversion from float to half-precision to invoke the half-precision sparse kernels as done in DGL or b) be forced to invoke float-based sparse kernels otherwise. The exponential has already performed one data conversion from half to float type of its input tensor.
PyTorch mix-precision stipulates that there is a real chance that half-precision is more likely to overflow when computing individual exponential as it can produce value in the range (0, INF), However, our studies for attention-based GNNs show that exponential operation cannot cause overflow because the model generates a non-positive number for attention coefficient ($e_{ij}$) (computed before Eq. 2). This observation is derived from DGL’s implementation of attention score [^1] https://github.com/dmlc/dgl/blob/master/python/dgl/backend/pytorch/sparse.py##L685, which applies an additive normalization. It first finds the maximum attention coefficient within a neighborhood, then each coefficient subtracts the maximum in that neighborhood, resulting in (-INF, 0] value range. Hence, its exponential is always in the range of (0, 1]. Therefore, the numerator of Eq. 2 can never overflow the half-precision.
This attention score implementation is slightly different than the original GAT model (gat18iclr, ), as the original GAT model has been shown to not converge in some datasets (featgraphsc20, ) even in float precision case.
General Trend. Exponential is not the only operator where mixed-precision automatically upgrades the incoming half-precision tensor to float before computing, but there are a range of operators that are used widely by multiple systems beyond attention-based GNNs that perform this upgrade (PyTorchAMP, ). The list includes common operations such as cross-entropy, log loss, softmax calculation, summation, etc. So almost all GNN should run into these operations, and would not be able to fully utilize half-precision kernels if the system is not designed accordingly. Hence, such system designs cannot be migrated to half-precision computation without performing excessive data-type conversion of tensors or invoking many float-based kernels thereby not providing much benefits in moving to mix-precision training.
3.1.3. Slower Half-Precision Sparse Kernels
We also observe that replacing float-based sparse kernels with half-precision kernels alone is not enough to improve system performance as the currently available half-precision based sparse kernels are slower than their float counterparts as shown earlier in Fig. 1(a) and Fig. 1(b). Specifically, DGL SDDMM, which is open-source, replaces float with the half-precision data-type directly without any system design change.
Based on this observation, we postulate that both the state-of-the-practice and the state-of-the-art for half-precision based GNNs differ from the corresponding dense deep learning model. The state-of-the-practice in half-precision based dense deep learning model focuses on the tensor core, which is well-suited to perform dense matrix multiplication. However, the tensor core is not suited due to the unstructured sparsity present in graph datasets used in GNNs. TC-GNN (wang2023tc, ) which was able to use tensor core in GNN does not perform better than float-based SpMM as shown by prior works (Graphpy, ; GNNBench, ).
Hence, half-precision GNNs must use half2 data-type, whose native support is available in GPUs, for better performance as shown earlier (ho2017exploiting, ) in the dense matrix multiplication before tensor core became prevalent. However, improving the state-of-the-practice by deploying the techniques from dense matrix multiplication (ho2017exploiting, ) alone would not be sufficient to achieve the accuracy and envisioned performance as we explore in this paper.
3.2. HalfGNN Overview
First of all, HalfGNN adopts the half2-based design and shows the challenges and solutions in adoption. We then show that the resultant design would continue to suffer from the overflow issue identified earlier as the root cause of the value overflow is reduction-related designs of float-based sparse kernels which are not suited directly for half-precision computation. We further show that reduction-related design continues to be the root cause of performance bottlenecks.
To this end, HalfGNN proposes half4 and half8 vector data-type to improve performance for SDDMM by minimizing the inter-thread communication cost which indirectly improves the instruction level parallelism to offer better data-load performance. We further propose discretized reduction scaling to protect against overflow in SpMM by exploiting the built-in degree-norm scaling. The proposed solution is also very well suited to provide workload balancing. We then introduce a staging buffer to remove atomic write, the root cause of performance bottleneck in half-precision SpMM.
The proposed solutions are generic and apply to edge-parallel and vertex-parallel sparse kernels. However, HalfGNN recommends an edge-parallel solution for the best performance.
4. Detailed Design
This section starts with adapting half2 to form a baseline and identifies that using half2 directly in a prior float-based state-of-the-art sparse kernel would not achieve the envisioned accuracy and performance goals without focusing on design around the reduction phase. We then discuss the SDDMM and SpMM design to overcome the issues.
4.1. Adapting Half2 Data-Type
Data-load phase implies bringing the graph and features to GPU registers for computation. This subsection show the challenges and techniques in adopting the half2 data-type.
4.1.1. Better Data-Load using Half2 Data Types
DGL SDDMM does not show any runtime improvement (Fig. 1(b)) as it merely replaces the float data-type with the half data-type. In such an approach, each thread brings one feature (2-byte) from GPU global memory to register for computation, hence a warp performs 64 bytes of data-load in case of full memory coalescing. In comparison, float-based SDDMM issues 128 bytes of data-load. Clearly, to fully utilize the memory transfer capability, it is mandatory to issue more data load to achieve full coalesced memory access. To this end, we adapt the half2 data type, provided natively by CUDA, which has not been used by prior GNN systems.
HalfGNN treats two contiguous half-precision features as one half2 feature, thus allowing a warp to issue 32 half2 values leading to issuing 128 bytes of data-load. Hence, HalfGNN achieves parity with float data-type for data-load performance in terms of size, but two times the number of features. This benefits SDDMM and SpMM both as the features, that we discuss next, are vectors.
We propose a split data-load phase to overcome the challenges associated with adopting half2.
Phase1: Explicitly Loading NZE and Edge-Features. They are explicitly loaded using the edge-parallel method and cached to GPU shared memory so that they can be read again very quickly when needed in the next phase. NZE is of integer data type while edge features, stored in row-major, are fetched sequentially. Hence we can still use vector data type to fetch vector features.
Kindly note that many prior SpMM works (gespmmsc2020, ; huang2021understanding, ; wang2021gnnadvisor, ) allocates a maximum of 32 edges to each warp, hence they cannot achieve better data-load performance for loading edge features, which is in half-precision now. To this end, we propose to allocate at least 64 edges to each warp, hence introducing one more design decision that can introduce better memory coalescing in this phase.
Edge-Feature Mirroring. We notice that only one of the halfs of the edge feature should do dot product with both halfs of half2 vertex features. To illustrate, Fig. 2 (d) for SpMM shows that only $W_{21}$ should be multiplied with {$X_{10}$, $X_{11}$}, but edge-feature data loading brings $W_{21}$ together with $W_{23}$ in our edge-centric design as half2 data-type. Therefore, performing dot product naively would also multiply $W_{23}$ with {$X_{10}$, $X_{11}$}, which is wrong.
Hence, after data loading edge features and before caching them in shared memory, each half is extracted from half2 edge features and is mirrored to create new half2 edge features which are then directly stored in the shared memory for caching. To illustrate, the first half2 word contains {$W_{21}$, $W_{21}$} (to be multiplied with {$X_{10}$, $X_{11}$}), and the next half2 contains {$W_{23}$, $W_{23}$} (to be multiplied with {$X_{30}$, $X_{31}$}). Performing this mirroring after data-load of edge features ensures that edge features can still use the half2 data-type for better data-load performance. Further, computation can be done without any issues or changes.
Phase2: Implicit Vertex Feature Loading. To load vertex-features, threads need to know the vertex ID of non-zero columns, which is already present in the shared memory thanks to phase 1. Vertex-features are stored in rowmajor order, and accessing 2 of them in a group is straightforward. As long as the feature-length is a multiple of 2, a simple type-casting of the features tensor to half2 allows us to use the half2 data type for data-loading. This allows 1 warp thread to carry out data load for 2 features at the same time, instead of 1. So, considering 32 threads in each warp, it can handle 64 features itself. So it doubles the number of features that can be loaded in parallel within each warp.
Feature Padding for Odd Feature Lengths. Kindly note that the feature-length is a configurable parameter in GNN, feature lengths such as 32, 64, etc. are common configurations. However, such lengths could be odd in the last layer of GNNs when the classification category itself is odd. E.g., Reddit has 41 classes to predict. In such cases, we simply use the next even number as class count without any issues.
Figure 3. Reduction in Edge-Parallel SpMM of HalfGNN: Showing 2 CTAs, each containing 4 warps. A warp handles 128 (configurable) edges.
Sub-Warps. We observe that multiple vertex-level features are loaded in a feature-parallel way within a warp. Hence, the threads being used are now half of what would be used in float-based sparse kernels. E.g., when the feature-length is 32, a whole warp processes an edge at any time in the case of float data type. However, using half2, we only need 16 threads of any warp to handle the same features. Hence, we process two non-zero columns at one time by the warp by dynamically creating two sub-warp within a warp. This allows us to use all the threads of the warp leading to better thread utilization. Indeed, we generalize this based on the feature-length, and hence the number of sub-warps within a warp depends on the feature length. E.g., in the case of 16 as feature length, we create 4 sub-warps.
4.1.2. Efficient Computation through Half2 Data-Type
Half arithmetic isn’t natively supported. If one relies on native operators, such as ‘+’ or ‘*’, the data is implicitly converted to float, and the operation is performed in float. The data conversion would be unnecessary. One can use CUDA intrinsic APIs for half computation that avoids this data-type conversion, but it only achieves the same throughput as the float. However, since we already perform data-load in half2, and CUDA intrinsic also provides native support for half2 arithmetic, we use half2 directly for dot product that performs multiplication of two halfs of half2 together. This approach doubles the computation arithmetic throughput.
4.2. SDDMM Reduction: More Vector Data-Types
SDDMM reduction happens across feature dimension that requires inter-thread communication for reducing the vector dot product to a single scalar. Inter-thread communication implies a memory barrier, and hence it does not allow the compiler to reorder memory instructions (e.g., data loads) across the memory barrier, and makes warp threads wait here until all threads finish their instruction issue. Hence, threads only issue one data-load if using half or half2, each thread loads only 2 half features. Further, a warp performs four rounds of inter-thread communication to reduce 32 features when using half2 data-type.
Thus the baseline version of SDDMM using half2 data type does not have an optimal design. Reliance on half2 alone is not sufficient to realize the full data-load potential of GPUs as we show in the evaluation that half2 cannot achieve the desired performance. HalfGNN proposes new optimizations to fully take advantage of half-precision in terms of two new vector data types, as discussed next.
Half4 and Half8. We propose, for the first time, two new vector data types, half4, and half8 similar to float2 and float4 respectively. The proposed half4 provides parity with float2 where 4 half-precision features are grouped as one half4 data point. Similarly, half8 treats 8 consecutive half-precision features as one half8 feature. However, there isn’t native support for these two data types, though CUDA supports float2 and float4 respectively. Hence, the proposed half4 internally uses float2 data-type to issue data-load, i.e. each warp can issue 256 bytes of data-load in a single data-load instruction (4×4$\times$4 × the number of features compared to float). Similarly, half8 internally relies on float4 data type to issue 512 bytes of data-load using a single data-load instruction (8×8$\times$8 × the number of features to float). Continuous memory locations of features can be accessed using a pointer data types of half, half2, half4, and half8 for read-write operation memory constraints are maintained. For example, if we have a pointer pointing to an array of data-type half, we can easily switch its pointer type from half to half2, half4, and half8 if the array size is a multiple of 2, 4, and 8 respectively. Thus, HalfGNN offers a range of vector data types to researchers.
For SDDMM, we use half8 data-type wherever possible. It then performs a dot product of vertex features of rows and column IDs of each NZE. Using half8, each thread loads 8 half features of columns, minimizes the inter-thread communication as half8 data-type requires only two rounds of inter-thread communication as only 4 threads are needed to handle 32 features compared to 4 rounds in half2-based SDDMM. Hence, by using half8 (i.e. 4 half2) for data-load, we observe an increase in instruction-level parallelism (ILP) compared to using half2 for feature load as each thread can now issue 4 half2 feature load before waiting at the memory barrier. Better ILP leads to better performance. The sub-warp concept( section 4.1.1) helps us utilize the idle threads better when using these vector types.
However, please note that, like the float2 and float4, half4 and half8 do not improve computation throughput than half2. Hence, even though we provide arithmetic on half4 and half8 data types, it internally relies on half2 arithmetic.
4.3. SpMM Reduction
For SpMM, one can also use half4 or half8, though it does not suffer from the issues that we outlined for SDDMM due to the absence of inter-thread communication. Specifically, reduction in SpMM is performed in the neighborhood dimension, so we need inter-thread communication only when the whole row is not assigned to a single warp. Indeed, Float-based SpMM has achieved workload balancing by dividing the whole neighborhood of any row to multiple warps. Fig. 3 shows the proposed edge-parallel HalfGNN SpMM, which we explain shortly. However, we observe that almost all state-of-the-art workload-balanced SpMM follow these rules:
  1. (1)
    Every edge is uniquely mapped to a single warp and consequently to a CTA, where its threads load vertex features in a feature-parallel way.
  2. (2)
    More than one edges are allocated to the same warp or CTA, whereby threads handle them in a loop. The consecutive edges follow the spatial ordering hence their row ID may be the same or in monotonically increasing order.
  3. (3)
    Edges distributed to consecutive CTA may share the same row ID if a row has large non-zero columns.
Rule 2 implies that a warp or CTA handles only a group of edges at any time, and is the main distinguishing factor between vertex-parallel and edge-parallel workload-balanced solutions. Specifically, in an edge-parallel design, a row-split may be observed within a warp for the latter as shown in Fig. 3. In a vertex-parallel design, rowID remains the same. For example, Huang et al (huang2021understanding, ) and GNNAdvisor (wang2021gnnadvisor, ) form a batch of 32 neighbors operated by each warp. Even vanilla vertex-parallel SpMM that offers no workload balancing, such as GE-SpMM (gespmmsc2020, ) has an implicit batching where each warp handles 32 neighbors of any vertex before processing the next 32 batches of neighbors for the same vertex in the next iteration, except the last iteration where it may process less than 32 neighbors depending on the vertex degree.
Irrespective, the design implies that workload-balanced solutions are easily amenable for discretized reduction scaling to protect against overflow, as discussed next.
4.3.1. Discretized Reduction Scaling
The proposed method is based on the observation on how the reduction is performed currently, and how can it be tailored to solve the problem. In effect, SpMM + degree-norm is treated as a custom SpMM kernel with ‘mean’ as a reduction operation, thus solving the problem for both GCN and GIN models.
We term the currently used reduction approaches as post-reduction scaling (degree-norm scaling applied once at the end of the reduction of the full neighborhood of vertices), performing exactly one division per vertex at the end of reduction. The other end of the spectrum would be to perform degree norm scaling before reduction, i.e., divide each dot product by degree and then perform reduction, pre-reduction scaling. This would perform exactly $|N(v)|$ count of division for each vertex, would avoid INF generation, but would introduce more arithmetic operations. While a few values might vanish (underflow) due to division at each dot product.
The proposed discretized reduction scaling shares the pros of both and the cons of none, as degree-norm scaling applies after reduction on a batch of neighbors for each vertex (rule 2). To illustrate, we perform the degree-norm scaling at the end of the reduction for each batch. Thus discretizing the reduction never risks getting INF as long as the the maximum number of neighbors in the batch is small. As one iteration of computation for a row handles a handful of neighbors, each iteration does produce a partially reduced value that never overflows on which the degree-norm scaling is performed. Thus, the proposed design changes are minimal and effective.
(3) $h_{u}=\phi\left((1+\epsilon)\cdot x_{u}+\sum_{v\in N_{v}}x_{v}\right)$
Additional Overflow in GIN. Discretized Scaled SpMM completely removes accuracy issues for GCN, but it is still not enough for GIN. In GIN aggregation, we observe 2 parts, one is the right representing unweighted SpMM being added to scaled source node features as shown in Eq. 3. The overflow issue does occur when adding self features for sources to SpMM output. We introduce a non-learnable parameter $\lambda$, to scale the aggregation down. Ideally, $\lambda=0.1$ worked fine for all our robust testing. So, we propose the following kind of aggregation for GIN in half-precision:
(4) $h_{u}=\phi\left((1+\epsilon)\cdot x_{u}+\lambda\cdot\left(\frac{1}{|N_{u}|}\sum_{v\in N_{u}}x_{v}\right)\right)$
4.3.2. Non-Atomic Writes
The discretized reduction also introduces conflicting write for which prior float-based works have relied on atomic write for the data-store phase, which is very costly for half2-based atomic writes. We first lay out when conflicting writes occur and how to resolve them: a) if one or more rows are contained exclusively within a warp, there is never any conflicting write, and hence no atomic write is needed. b) If one or more rows are contained exclusively within a CTA, then an intra-CTA communication can resolve those conflicting writes locally within the CTA to generate non-conflicting writes. We discuss this soon. c) Only when a row extends to more than one CTA that conflicting writes cannot be resolved. To resolve this case, we work with point (b) above and deploy a staging buffer to direct the conflicting write. The staging buffer is then re-read in a separate follow-up kernel to write to the output tensor. Fig. 3 shows the details, discussion follows next.
Conflict Write Handling. The thread-local reduced value is written non-atomically to output tensor directly in case of row-split as shown for the yellow row in Fig. 3. The reduced values at warp-end (non-row-split case) are the ones that are resolved first locally within a CTA by using shared memory. Kindly note that current libraries, such as Nvidia’s CUB offer intra-CTA communication only for scalar dot products for reduction, and are insufficient for exchanging vector dot products.
Figure 4. HalfGNN achieves the same accuracy as float-based DGL, confirming the applicability of the proposed techniques. Only labeled datasets are used to show accuracy.
We propose a new design for intra-CTA communication, which also handles the case of sub-warp (section 4.1.1). In such a case, not all threads need to communicate with every other thread; the communication happens between a thread of a sub-warp to the same thread ID of all other sub-warps of the CTA so that the same feature gets reduced. Hence, if there are 8 sub-warps within a CTA, we only perform three rounds of communications within the CTA. The proposed library relies on cached row ID stored in the shared memory to know if the dot products stored in a sub-warp should be merged to the dot products of another sub-warp.
The resolved writes are written directly to the output tensor. The unresolved writes are the ones that can only be resolved by looking into the writes by the next CTAs. We treat the last (partial) row assigned to this CTA as the conflicting write. This generates just one final conflicting write of size feature-length, which is the output corresponding to the row ID of the last edge(s) assigned to this CTA and are written to the staging buffer. A follow-up kernel is written to handle this buffer to perform the final write.
Staging Buffer & Writing Heuristic. This buffer is where we store the carry-outs is of size $|CTA|\times 4$ bytes. The number of total destination vertices is $|CTA|$ and the memory requirement is $|CTA|\times 4$ bytes. This combined data structure requires little memory, and we save a lot by using half-precision for state tensors already.
4.3.3. Generality of the Solution
Our non-atomic design principle involves intra-CTA, & inter-CTA conflict resolution, with 2-stage kernel design. This applies to edge-parallel design as HalfGNN itself is edge-parallel. Not only that, this applies to even vertex-parallel designs. We apply the vectorization design and non-atomic write to Huang et al  (huang2021understanding, ) and gain 1.79×1.79$\times$1.79 × speedup on average for SpMM. This shows the proposed optimizations can massively improve prior works.
4.4. Shadow APIs
Figure 5. HalfGNN requires much less memory to train compared to DGL float. (Lower is better)
In mixed-precision training, DL frameworks invoke float versions of many operations (PyTorchAMP, ), e.g. exponential (section 3.1.2). However, if we are sure that such operations would produce the output within the numerical range of half-precision, we advocate using the half-precision version. This guarantees an optimal number of kernel calls are half-precision and the least number of float conversions for maximum performance.
However, DL frameworks, such as PyTorch, do not automatically invoke such operation in half-precision (PyTorchAMP, ). To this end, we provide similar APIs, called shadow API that guarantee that the half-precision version would be invoked if the input tensor is of half data-type, and the float version would be invoked if the input tensor is of float type. Hence, one can always use the proposed shadow API if we are sure that overflow would not incur in half-precision.
5. Evaluations
The following experiments aim to comprehensively evaluate the proposed optimizations and perform a comparative study compared to prior GNN works. All the experiments are conducted in NVIDIA A100 GPU which has 40GB of GPU memory. We conducted our experiments on 16 datasets mentioned in Table 1. Only 5 datasets (Cora, Pubmed, Citeseer, Ogb-product, and Reddit) are labeled where accuracy is measured, whose input feature lengths are shown. Others are used for evaluating runtime, and memory usage only, hence they contain generated features.
Table 1. Graph datasets. * denotes labeled dataset. F = Input feature-length, C = Prediction categories
Graph Dataset $|V|$ $|E|$ $|F|$ $|C|$
Cora (G1)* 2,708 10,858 1,433 7
Citeseer (G2)* 3,327 9,104 3,703 6
PubMed (G3)* 19,717 88,648 500 3
Amazon (G4) 400,727 6,400,880 150 7
Wiki-Talk (G5) 2,394,385 10,042,820 150 7
RoadNet-CA (G6) 1,971,279 11,066,420 150 7
Web-BerkStand (G7) 685,230 15,201,173 150 7
As-Skitter (G8) 1,696,415 22,190,596 150 7
Cit-Patent (G9) 3,774,768 33,037,894 150 7
Sx-stackoverflow (G10) 2,601,977 95,806,532 150 7
Kron-21 (G11) 2,097,152 67,108,864 150 7
Hollywood09 (G12) 1,069,127 112,613,308 150 7
Ogb-product (G13)* 2,449,029 123,718,280 100 47
LiveJournal (G14) 4,847,571 137,987,546 150 7
Reddit (G15)* 232,965 114,848,857 602 41
Orkut (G16) 3,072,627 234,370,166 150 7
5.1. GNN Training
This sub-section evaluates training results, showing accuracy, runtime, and memory usage. We focus on GCN, GAT, and GIN models to show the overall effectiveness of HalfGNN. These models contain intermediate feature lengths of 64. The intermediate feature length is important as the sparse kernels usually have this feature length during computation.
5.1.1. Training Accuracy
As we highlighted earlier in Fig. 1(c), the accuracy of DGL for half-precision is much lower than the float-based models, which questioned whether half-precision is a viable option for GNN training. Hence, we use DGL single-precision as the baseline in Fig. 4 to show that HalfGNN achieves similar accuracy to the baseline. The difference in accuracy is less than 0.3% for all except pubmed in GIN for which it is less than 1.0% improvement. Firstly, these differences are mostly due to random weight initialization, and changes by this margin if re-run. Secondly, we have made a minor adjustment to protect the additional overflow as discussed in section 4.3.1. This implies that half-precision is a viable option in place of float without compromising the GNN accuracy.
The reason for the similar accuracy in HalfGNN as float-based DGL is protecting the overflow through the proposed discretized reduction scaling in SpMM. Protection for SpMM is possible only at the sparse kernel level as frameworks themselves cannot protect against such overflow. To confirm this result, we replaced the discretized reduction with the usual reduction so that overflow would not be protected. This system did result in DGL-half like abnormal accuracy.
5.1.2. Training Runtime
Fig. 6 shows the runtime speedup of training time of GCN, GAT, and GIN using HalfGNN over DGL. All GNNs were trained for 400 epochs. Kindly recall that Fig. 1(a) and Fig. 1(b) have shown that half-precision based DGL kernels are slower than float-based DGL, while showing abnormal accuracy. Hence, we picked the faster (and correctly executed) baseline here. Compared to DGL float HalfGNN achieves 1.85×,3.55×1.85$\times$,3.55$\times$1.85 × , 3.55 ×, and 1.78×1.78$\times$1.78 × speedup on average for GCN, GAT, and GIN respectively, providing around an average of 2.30×2.30$\times$2.30 × speedup over float-based DGL
Figure 6. HalfGNN training runtime speedup over DGL-float for feature size 64. DGL runs out of memory for GAT G16. (log scale, higher is better)
Thus, the results so far show that half-based GNN is not only feasible(section 5.1.1) but is also significantly faster to train. The reason for the faster training time is the better sparse kernels of HalfGNN compared to prior GNN works, specifically introducing half-precision specific performance optimization to the sparse kernels. We show the runtime comparison of these kernels next by focusing on evaluating the impact of the proposed optimizations of HalfGNN.
5.1.3. Memory Consumption
HalfGNN also achieve 2.67$\times$ less memory consumption on average than float-based DGL considering all 3 GNNs as plotted in Fig. 5. While it is clear that half-precision would improve memory usage, one must note that the advantage is not purely due to moving to half-precision as DGL consumes excessive memory due to its design choices and has framework-memory overhead as shown recently (GNNBench, ).
5.2. Kernel Evaluation
In this section, we focus on individual kernel performances compared to established DGL/Cusparse kernel performance and compare with our results.
Figure 7. HalfGNN achieves 22.89×22.89$\times$22.89 × average speedup over DGL half-precision SpMM (Cusparse), and 7.12×7.12$\times$7.12 × average speedup over DGL half-precision SDDMM. Speedup over 64 has been clipped. (Log Scale, higher is better) )
5.2.1. SpMM Runtime
DGL uses Cusparse for SpMMve. We measure the runtime for feature sizes 32 and 64. We compared our runtime against the Cusparse half-precision. Fig. 7 shows the comparative speedup. We can observe that the speedup is massive for all datasets compared to the Cusparse. Furthermore, these speedups are consistent over feature sizes 32, and 64. HalfGNN SpMMve achieves over 64×64$\times$64 × runtime speedup in some cases, while achieving on average 22.89×22.89$\times$22.89 ×.
These speedups are observed due to two reasons: a) better design of half-precision kernel by HalfGNN which we have explored separately in section 5.3, and b) huge performance degradation by Cusparse half-precision library compared to its own float-based SpMM as discussed earlier in Fig. 1(a). Indeed, the huge speedup is due to the latter, as our additional measurement shows that HalfGNN achieves a more realistic 2.52×2.52$\times$2.52 × speedup over Cusparse-float.
SpMMv shares the same design principles and optimizations as SpMMve, whereas Cusparse does not provide this custom SpMM. Hence, we have not plotted it separately. This faster SpMM helps improve the training time for GCN, GAT, and GIN.
5.2.2. SDDMM Runtime
Fig. 7 shows that HalfGNN gets up to 12×12$\times$12 × speedup, while most datasets get over 6×6$\times$6 × speedup compared to the half-precision DGL SDDMM. The average speedup is around 7.12×7.12$\times$7.12 ×. The reason behind such massive speedup is that DGL half-precision kernels individually do not provide any runtime advantage over their float counterparts as it just replaces float data-type with half without even using half2 or any other optimizations, as shown earlier in Fig. 1(b). On the other hand, HalfGNN adopts half8 vector data-type which provides a huge performance boost in data-load time, increased instruction-level parallelism, and fewer rounds of inter-thread communication within a warp.
The speedup is huge for half-precision individual kernels of SpMM and SDDMM, yet we have shown only 2.30$\times$ speedup in training in Fig. 6. The realistic speedup is because we plotted float-based DGL in Fig. 6 as it trains faster than half-precision DGL, and trains correctly.
5.3. Micro-Benchmarking
In this section, we perform many micro-benchmarking to measure the impact of our proposed optimization choices.
5.3.1. SDDMM: Half8 vs Half2
We now show that usage of half8 is more efficient than using only half2 for SDDMM. Fig. 8 shows that using half8 can provide runtime speedup up to almost 3×3$\times$3 ×, and achieves an average speedup of 1.67×1.67$\times$1.67 × for feature sizes $32$ and $64$. Using half8 reduces the number of required inter-thread communications for the reduction in SDDMM and allows to issue load of four half2 features compared to just one , which results in better data-load performance due to increased instruction-level parallelism.
Figure 8. Using half8 achieves 1.67×1.67$\times$1.67 × speedup over half2 based SDDMM (Higher is better)
5.3.2. SpMM: Atomic vs. Non-atomic Writes
Fig.  9 shows the runtime speedup achieved when HalfGNN removes the atomic write through intra-CTA communication, staging buffer, and follow-up kernel compared to a version that uses atomic write, keeping everything as is for edge-parallel design. This is true not only for Reddit and Ogb-product datasets but also for all the other datasets. Atomic implementations are more costly for half-precision than float data-type. A large part of our design for SpMMve revolved around how to get rid of atomic write. We believe, HalfGNN is the first-ever implementation, demonstrating the performance gain using non-atomic vectorized kernels in half-precision.
Figure 9. Speedup when atomic ops removed from HalfGNN SpMM
5.3.3. Huang-float vs Huang-Half2
. We chose Huang et al (huang2021understanding, ), a state-of-the-art workload-balanced vertex-parallel work, to showcase the effectiveness of our optimization proposed in HalfGNN. Huang et al SpMM groups the neighborhood of every vertex in the batch of 32, which is assigned to one warp, and hence uses atomic write to resolve conflicting writes in its float-version. We adopt it to half-precision, where we use half2, and remove the atomic write using staging buffer and follow-up kernel. However, the full application of HalfGNN SpMM does not apply to this approach specifically the edge-parallel approach. Further, a warp performs data loading edge features of 64 bytes and not 128 bytes due to batching only 32 neighbors thereby compromising the memory coalescing. Lastly, the data-load phase of edge features does not always start at even offset, leading to some compromise on the usage of half2 data type.
Figure 10. Huang-half2 gains 1.79×1.79$\times$1.79 × speedup over Huang-float.
Fig. 10 shows their runtime comparison showing that the adopted half-precision version achieves around 1.79×1.79$\times$1.79 × speedup on average. This is massive, as we achieve this without changing any fundamental nature of Huang et. al, without applying even the full range of proposed optimizations. This demonstrates the generality of our proposed optimization. Even though HalfGNN is an edge-parallel system, its design can massively improve performance for vertex-parallel systems also.
6. Conclusion
We studied GNN training in half-precision for the first time. We identified major issues that prevent existing half-precision systems from training GNN correctly. We provided generic solutions for that to explore half-precision training. Furthermore, we introduced novel techniques that can take advantage of half-precision to achieve state-of-the-art system performance for SpMM, SDDMM, and overall GNN training. We provided a system that integrates all these.